# Ultra-Sparse Memory Network

### v1

@@ Abstract
It is widely acknowledged that the performance of Transformer models is exponentially related to their number of parameters and computational complexity. While approaches like Mixture of Experts (MoE) decouple parameter count from computational complexity, they still face challenges in inference due to high memory access costs. This work introduces UltraMem, incorporating large-scale, ultra-sparse memory layer to address these limitations. Our approach significantly reduces inference latency while maintaining model performance. We also investigate the scaling laws of this new architecture, demonstrating that it not only exhibits favorable scaling properties but outperforms traditional models. In our experiments, we train networks with up to 20 million memory slots. The results show that our method achieves state-of-the-art inference speed and model performance within a given computational budget.
@@ 3 UltraMem
Product-key memory layers scale up the memory size with $N>10^{6}$, while activating only a few values with top-$m$ scores. Here, $m$ is a hyper-parameter controlling sparsity. Though values are sparsely accessed, the keys, which are as large as values, must be fully computed to obtain scores before top-$m$ activation following equation 1. To alleviate the computation complexity for keys, product keys are proposed. It utilizes a 2-D logical address (see Figure 2(b)), typically a $n\times n$ grid where $n=\sqrt{N}$, for memory value retrieval. Specifically, a 2-D logical address $(i,j)$ is used to index memory value at physical address $n\times i+j$. With such strategy, logical scores are then represented as a matrix, which is further decomposed as an addition of row and column scores:
@@ 3.2 Structure improvements
Figure 4: Flow of Tucker Decomposed Query-Key Retrieval, here $r=2$. “fetch” refers to retrieving score based on given index.
@@
Figure 5: Flow of Implicit Value Expansion, here $E=4$, $m=16$.
@@ 5.1 Setup
Training details. We used a standard pre-norm transformer (Xiong et al., 2020) with rotary embeddings (Su et al., 2024). Our dense models are built with 151M [^2] Parameter counts in this paper exclude tokenizer vocabulary embedding and prediction head parameters., 680M, 1.6B, and 6.5B parameters. For sparse models, including UltraMem and MoE, we expand the sparse parameters twelve fold from the 151M, 680M, and 1.6B dense models. In MoE models, two experts are activated per token (Jiang et al., 2024), and we use a standard balance loss (Fedus et al., 2022) with a weight of 0.01 to ensure an even selection of all experts, thereby establishing MoE as a strong baseline. We slightly increase the width of MoE’s experts to align the computational and parameter costs with those of the UltraMem. In UltraMem models, auxiliary loss weight $\alpha=0.001$, margin $\tau=0.15$, the learning rate for values is ten times that of other parameters and linearly decays to equal the other parameters by the end of training. For details on model structure and hyperparameters, see Appendix E. For details on the optimizations made for large-scale training, please see the Appendix C,D.
@@ 5.2 Evaluation on Language Modeling Datasets
We evaluate models of various sizes, the results are shown in Table 1333This table only includes evaluation results where the metrics have steadily increased with training. For all results, see the Table 7 in Appendix., where FLOPs is the computation cost of single token, the curves showing changes over the course of training are provided in the Figure 11 in Appendix. We observe that as the model capacity increases, UltraMem can outperform MoE with the same parameter and computation. On the 1.6B dense model, an UltraMem model with 12x the parameters can match the performance of a 6.5B dense model.
Table 1: Performance metrics of various models
@@ 5.3 Value number and top-$m$
As sparse parameters increase, we observed differences in inference speeds between UltraMem and MoE, as depicted in Figure 6(c). For UltraMem, the inference time remains largely stable even as sparse parameters exponentially grow, provided that the activation parameters (top-m) are constant. In contrast, MoE’s inference time escalates significantly under analogous conditions. Referring to Figure 1(b), with smaller batch sizes, the inference speed of MoE deteriorates even further compared to UltraMem.
@@ 5.4 Ablation
Training Validation Dense Sparse FLOPs
Loss $\downarrow$ Loss $\downarrow$ Params(M) Params(G) (M)
@@
Among other changes, sharing the query helps cut computational costs with a minor trade-off in performance. Normalizing the query/key greatly reduces spikes in training perplexity and enhances training stability, as shown in Figure 10(a). Improved initialization prevents score and output variance explosions in the early to middle training stages, detailed in Figure 10(b) and (c). Additionally, employing convolution further limits the variance divergence in UltraMem outputs(Figure 10.(c)).
@@ Appendix E Experiment setting
Table 4 displays common hyper-parameter settings for all experiments. “LR” stands for Learning Rate, corresponding to the values 6e-4, 2.5e-4, 2e-4, and 1.2e-4 for dense models with sizes 151M, 680M, 1.6B, and 6.5B, respectively (Brown, 2020). Regarding the insertion of UltraMem, for UltraMem-151M, it’s 3:5/6:8/9:11, where 3:5 indicates that UltraMem input is taken from layer 3 and inserted back into the output of layer 5, and so on. For UltraMem-680M, it’s 3:7/8:12/13:17/18:22. For UltraMem-1.6B, it’s 3:7/8:12/13:17/18:22/23:27/28:32. The settings for UltraMem and MoE models align with their dense counterparts based on dense parameter size. Table 6 shows the model parameter setting used in scaling experiments. What’s more, the common setting for UltraMem is shown in Table 5.
@@
Model Hidden Dim Inner Dim Attn Head Layer Top-m Expert Kdim Knum Ultra Mem Layer Param (B) FLOPs (G) Dense-151M 1024 4096 16 12 - - - - - 0.15 0.30 Dense-680M 1536 6144 16 24 - - - - - 0.68 1.36 Dense-1.6B 2048 8192 16 32 - - - - - 1.61 3.21 Dense-6.5B 4096 16384 32 32 - - - - - 6.44 12.88 MoE-151M-2in32 1024 2528 16 12 2 32 - - - 2.04 0.35 MoE-680M-2in33 1536 3584 16 24 2 33 - - - 8.95 1.50 MoE-1.6B-2in34 2048 4672 16 32 2 34 - - - 21.36 3.52 UltraMem-151M-x10 2048 8192 16 32 42x2 - 256 1024 3 1.71 0.35 UltraMem-151M-x12 1024 4096 16 12 16x2 - 256 1100 3 2.03 0.35 UltraMem-680M-x12 1536 6144 16 24 35x2 - 384 1632 4 8.93 1.49 UltraMem-1.6B-x12 2048 8192 16 32 42x2 - 448 1792 6 21.41 3.50
Table 6: Model parameter setting. Top-$m$ means chosen expert number in MoE, means chosen value number times head number in UltraMem. Kdim means the key dimension in UltraMem. Knum means the number of keys, Knum2 is the number of values.

### v2

@@ Abstract
It is widely acknowledged that the performance of Transformer models is logarithmically related to their number of parameters and computational complexity. While approaches like Mixture of Experts (MoE) decouple parameter count from computational complexity, they still face challenges in inference due to high memory access costs. This work introduces UltraMem, incorporating large-scale, ultra-sparse memory layer to address these limitations. Our approach significantly reduces inference latency while maintaining model performance. We also investigate the scaling laws of this new architecture, demonstrating that it not only exhibits favorable scaling properties but outperforms MoE. In experiments, the largest UltraMem we train has 20 million memory slots. The results show that our method achieves state-of-the-art inference speed and model performance within a given computational budget, paving the way for billions of slots or experts.
@@ 3 UltraMem
Product-key memory layers scale up the memory size with $N>10^{6}$, while activating only a few values with top-$m$ scores. Here, $m$ is a hyper-parameter controlling sparsity. Though values are sparsely accessed, the keys, which are as large as values, must be fully computed to obtain scores before top-$m$ activation following equation 1. To alleviate the computation complexity for keys, product keys are proposed. Borrowing the idea of Product Quantization, it utilizes a 2-D logical address (see Figure 2(b)), typically a $n\times n$ grid where $n=\sqrt{N}$, for memory value retrieval. Specifically, a 2-D logical address $(i,j)$ is used to index memory value at physical address $n\times i+j$. With such strategy, logical scores are then represented as a matrix, which is further decomposed as an addition of row and column scores:
@@ 3.2 Structure improvements
Figure 4: Flow of Tucker Decomposed Query-Key Retrieval (TDQKR), here $r=2$. The term “fetch” refers to the action of retrieving scores based on a given index (corresponding to “torch.gather”). TDQKR replacing Product Quantization, serves as a more precise retrieval module for recalling value indices in UltraMem. Each step of the TDQKR process is meticulously referenced within the main text for understanding.
@@
Figure 5: Flow of Implicit Value Expansion (IVE), here $E=4$, $m=16$. IVE reduces memory access and scales up memory size by expanding the memory table virtually. Each virtual block is a reparameterization of the physical memory table. Every virtual memory address corresponds to a physical memory address and a projector index. The weighted sum pooling is grouped by the virtual blocks, followed by a linear layer to produce the final output.
@@ 5.1 Setup
Training details. We used a standard pre-norm transformer (Xiong et al., 2020) with rotary embeddings (Su et al., 2024) for our dense models, which have 151M, 680M, 1.6B, and 6.5B parameters [^2] Excludes tokenizer vocabulary embedding and prediction head parameters.. For sparse models, including UltraMem, PKM and MoE, we expand the sparse parameters twelvefold from the 151M, 680M, and 1.6B dense models. In MoE models, two experts are activated per token (Jiang et al., 2024), using a balance loss (Fedus et al., 2022) weight of 0.01 to ensure even expert selection. We slightly increased the width of MoE’s experts to match UltraMem’s computational and parameter costs. In UltraMem models, the auxiliary loss weight is $\alpha=0.001$ and margin $\tau=0.15$. The learning rate for values is ten times other parameters and decays linearly. For model structure and hyperparameters details, see Appendix E, and for large-scale training optimizations, see Appendix CD.
@@ 5.2 Evaluation on Language Modeling Datasets
We evaluate models of various sizes, the results are shown in Table 1333This table only includes evaluation results where the metrics have steadily increased with training. For all results, see the Table 7 in Appendix., where FLOPs is the computation cost of single token, the curves showing changes over the course of training are provided in the Figure 11 in Appendix. We observe that as the model capacity increases, UltraMem can outperform PKM and MoE with the same parameter and computation. On the 1.6B dense model, an UltraMem model with 12x the parameters can match the performance of a 6.5B dense model.
Table 1: Performance metrics of various models.
@@ 5.3 Value number and top-$m$
As sparse parameters increase, Figure 6(c) shows that UltraMem maintains stable inference time despite exponential growth in parameters, as long as activated parameters (top-m) stay constant. In contrast, MoE’s inference time rises significantly under similar conditions. Additionally, Figure 1(b) demonstrates that with smaller batch sizes, MoE’s inference speed deteriorates even more compared to UltraMem.
@@ 5.4 Ablation
Train Loss $\downarrow$ Valid. Loss $\downarrow$ Dense Param.(M) Sparse Param.(G) FLOPs (M)
@@
Among other changes, sharing the query helps cut computational costs with a minor trade-off in performance. Normalizing the query/key greatly reduces spikes in training perplexity and enhances training stability, as shown in Figure 10.(a). Improved initialization prevents score and output variance explosions in the early to middle training stages, detailed in Figure 10.(b) and (c). Additionally, employing convolution further limits the variance divergence in UltraMem outputs(Figure 10.(c)). The above results are based on incremental ablation. Results from independent ablation can be found in the Table 8 in Appendix, and they align with our expectations.
@@ 6 Conclusion
Acknowledgments
We extend our deepest gratitude to Pingshuo Ma and Wenda Liu for their invaluable assistance in optimizing the early stage training of large-scale UltraMem. We also appreciate the inference optimization work for the UltraMem carried out by Siyan Chen, as well as Fan Xia’s efforts in assessing the inference speed of MoE.
@@ Appendix E Experiment setting
Table 4 displays common hyper-parameter settings for all experiments. “LR” stands for Learning Rate, corresponding to the values 6e-4, 2.5e-4, 2e-4, and 1.2e-4 for dense models with sizes 151M, 680M, 1.6B, and 6.5B, respectively (Brown, 2020). Regarding the insertion of UltraMem and PKM, for UltraMem-151M, it’s 3:5/6:8/9:11, where 3:5 indicates that UltraMem input is taken from layer 3 and inserted back into the output of layer 5, and so on. For UltraMem-680M, it’s 3:7/8:12/13:17/18:22. For UltraMem-1.6B, it’s 3:7/8:12/13:17/18:22/23:27/28:32. For PKM-151M, it’s 6:6. For PKM-680M, it’s 12:12. For PKM-1.6B, it’s 16:16. The settings for UltraMem and MoE models align with their dense counterparts based on dense parameter size. Table 6 shows the model parameter setting used in scaling experiments. What’s more, the common setting for UltraMem is shown in Table 5.
@@
Model Hidden Dim Inner Dim Attn Head Layer Top-m Expert Kdim Knum Mem Layer Param (B) FLOPs (G) Dense-151M 1024 4096 16 12 - - - - - 0.15 0.30 Dense-680M 1536 6144 16 24 - - - - - 0.68 1.36 Dense-1.6B 2048 8192 16 32 - - - - - 1.61 3.21 Dense-6.5B 4096 16384 32 32 - - - - - 6.44 12.88 MoE-151M-2in32 1024 2528 16 12 2 32 - - - 2.04 0.35 MoE-680M-2in33 1536 3584 16 24 2 33 - - - 8.95 1.50 MoE-1.6B-2in34 2048 4672 16 32 2 34 - - - 21.36 3.52 PKM-151M-x12 1024 4096 16 12 16x6 - 512 1347 1 2.04 0.35 PKM-680M-x12 1536 6144 16 24 35x8 - 768 2308 1 8.95 1.50 PKM-1.6B-x12 2048 8192 16 32 42x12 - 896 1792 1 21.44 3.52 UltraMem-151M-x10 2048 8192 16 32 42x2 - 256 1024 3 1.71 0.35 UltraMem-151M-x12 1024 4096 16 12 16x2 - 256 1100 3 2.03 0.35 UltraMem-680M-x12 1536 6144 16 24 35x2 - 384 1632 4 8.93 1.49 UltraMem-1.6B-x12 2048 8192 16 32 42x2 - 448 1792 6 21.41 3.50
Table 6: Model parameter setting. Top-$m$ means chosen expert number in MoE, means chosen value number times head number in PKM and UltraMem. Kdim means the key dimension in PKM and UltraMem. Knum means the number of keys, Knum2 is the number of values.
@@ Appendix F More Experiment results
Table 8: Independent ablation study of model improvements
Train Loss $\downarrow$ Valid. Loss $\downarrow$
PKM-151M-x10 2.604 2.828
+ rm softmax -0.034 -0.006
+ half vdim+proj -0.027 -0.02
+ share query -0.003 -0.002
+ split big mem -0.003 -0.005
+ query/key LN -0.002 +0.003
+ IVE -0.025 -0.023
+ TDQKR -0.003 -0.007
+ TDQKR + MCS -0.02 -0.009
+ value lr decay -0.017 -0.007
+ query conv -0.005 -0.001